# A Survey of Protocol Fuzzing

### v2

@@ 5.1.2. Execution Cost Reduction
1. Reset Time Selection. The reset time selection has a significant influence on the performance of fuzzing. Early reset may cause the target to terminate when it is still doing some tasks that may be vulnerable, and late reset may affect the efficiency of the test. A common approach is to set a fixed time interval before resetting the execution. For example, AFLNET (Pham et al., 2020) allows the user to manually configure the time delays before restarting the PUT. However, this approach is relatively coarse-grained, and it is hard to determine an appropriate time interval. In order to precisely control when to reset the execution, some works use program analysis to find the location that indicates the end of the execution of an iteration, and instrument the target program to terminate at these code locations (Andronidis and Cadar, 2022; Schumilo et al., 2022; Krupp et al., 2022; Qin et al., 2023). For example, AMPFuzz (Krupp et al., 2022) performs static analysis and injects termination calls to the code branches that do not contain message-sending APIs. In addition, some works choose not to perform an execution reset after each fuzzing iteration for performance boost. For example, Charon (Zuo et al., 2022) leverage a program status inferring module to infer the time point at which the PUT has finished processing the packet, thereby detecting the coverage of specific inputs and avoid the need to repeatedly restart the PUT to collect feedback. Similarly, SGFuzz (Ba et al., 2022) doesn’t restart the PUT in every iteration. However, it performs a post-analysis to eliminate the uon-deterministic. Specifically, it collects all the inputs on which the PUT has been executed and minimizes the input list to a minimal message sequence that can trigger the bug.
@@
2. Execution State Reset. Execution state reset is responsible for resetting the context of the running PUT process to a specified state, including the data in registers and memory, etc. Existing execution state reset mechanisms can be divided into three categories, namely message-based reset, process restart, and snapshot & recovery.
@@
Message-based reset (labeled as “MR” in Table. 5 column 6) operates by dispatching a specific type of message that compels the PUT to terminate the ongoing session and revert to its initial state (Vanhoef, 2017; Fiterau-Brostean et al., 2020; Bytes et al., 2022). For instance, when fuzzing the Wi-Fi Access Point (AP), WiFuzz use a deauthentication message to reset its state (Vanhoef, 2017). Message-based reset is easy-to-use, but it only supports a limited set of protocols as not every protocol is designed with a reset message. Moreover, it can only reset the explicit protocol state of the PUT, but the implicit state of the test target, such as the global variables, and the memory allocated but never freed, cannot be reset with this approach.
@@
Another commonly used approach to reset the execution state is to kill the target process and restart (labeled as “ProcR” in Table. 5 column 6) (Chen et al., 2018; Luo et al., 2020; Fang et al., 2021). However, it is a relatively heavy operation for fuzzing, as the restart of the program involves multiple expensive pre-processing steps, such as loading the program into memory, dynamic linking etc., resulting in inefficiencies.
@@
The snapshot & recovery mechanism has been integrated into fuzzing. This approach involves checkpointing PUT at a specific runtime state and then resetting it back to that checkpoint after each fuzzing iteration. This method effectively bypasses the repeated execution of resource-intensive initialization operations, thereby enhancing fuzzing efficiency. Protocol fuzzing, in particular, derives significant benefits from snapshot technology. Protocols are predominantly stateful, implying that the input often comprises multiple prefix messages that guide the PUT to a designated state before introducing the crafted message. It’s common for testcases to share the same prefix message sequences, especially when a specific state requires repeated exploration. Implementing snapshot technology in the protocol fuzzing process eliminates redundant executions associated with parsing these shared packet sequences, thereby markedly boosting fuzzing efficiency. The snapshot methodologies employed in current protocol fuzzing research can be broadly categorized into two types: process-level snapshots and virtual machine-level snapshots.
@@
Process-Level snapshot mechanisms (labeled as “PSR” in Table. 5 column 6) rely on system call capabilities provided by the operating system to realize their functionality. Generally, based on the APIs used, existing methods can be categorized into two types: fork-based and ptrace-based. Fork-based snapshot mechanisms are widely used in several well-known general-purpose fuzzers, including AFL(Zalewski, 2015). Specifically, AFL inserts a piece of fork-server code into the PUT program binary, which is executed before the function. Following a signal from the AFL fuzzing side, the fork-server generates a child process via the function, and this child process continues with the function. Since the fork-server has already loaded all kinds of resources, each child process only needs to execute the main function’s code, thereby bypassing the costly pre-processing steps and enhancing efficiency. This mechanism has been adopted by many protocol fuzzers for state resetting (Luo et al., 2019; Pham et al., 2020; Natella, 2022; Maier et al., 2020; Li et al., 2022a; Andronidis and Cadar, 2022; Qin et al., 2023; Fu et al., 2023). Additionally, some works have extended the original fork-server mechanism in AFL to allow for conditional multiple initializations at different code points, enabling the fuzzer to conveniently switch between various states of the protocol and thus boosting the fuzzing process (Chen et al., 2019; Andronidis and Cadar, 2022). ptrace-based snapshot mechanisms, such as CRIU and DMTCP, leverage debugging API to collect all the process context information and save it as image files (Li et al., 2022a). In the restoration process, these snapshot mechanisms read the dumped image files and recreate the process using syscalls such as or . Unlike the fork-based snapshot, which requires predetermined snapshot conditions (i.e., the location of the fork-server call) before execution, ptrace-based snapshots can checkpoint at any state during runtime.
@@
Virtual-Machine-Level snapshot mechanisms (labeled as “VMSR” in Table. 5 column 6) utilize the capabilities of virtual machine hypervisors to capture snapshots of the entire virtual machine at a specific time point, typically facilitated through a hypercall (Schumilo et al., 2022; Shi et al., 2019). When hypercalls are invoked, the program running within the virtual machine exits the VM context and transfers control to the hypervisor. Although the hypervisor-based approach is user-friendly, requiring no instrumentation, it is somewhat less efficient and more space-consuming due to its large granularity. To enhance the practicality of using virtual machine-level snapshots in protocol fuzzing, Nyx-net (Schumilo et al., 2022) implements an incremental snapshot approach to reduce the overhead associated with creating and removing snapshots. Specifically, Nyx-net establishes a root snapshot in a pristine state, and each execution iteration commences from this root snapshot. In subsequent fuzzing iterations, Nyx-net generates incremental snapshots based on the root snapshot following the execution of an input message. Consequently, Nyx-net has a great performance boost on the testcases that share the same prefix message sequences.
@@
3. Execution Environment Reset. The reset of the execution environment primarily involves resetting the filesystem or database that may be affected by the PUT. Many fuzzers require users to provide a cleanup script to revert all changes (Pham et al., 2020; Qin et al., 2023; Fu et al., 2023; Li et al., 2022a; Natella, 2022), necessitating substantial manual effort to analyze the PUT’s potential impact on the external environment. To address this issue, Snapfuzz (Andronidis and Cadar, 2022) leverages a custom in-memory filesystem, wherein modifications are automatically discarded after the completion of a fuzzing iteration. Furthermore, the hypervisor-based snapshot mechanism. Besides, the hypervisor-based snapshot mechanism (VMSR in Table. 5 column 7) (Schumilo et al., 2022), which captures the state of the entire virtual machine, can reset both the execution state and the environment simultaneously.
@@ 5.2. Runtime Information Extraction
Software-Based methods leverage the capability of the software execution environment, e.g., compiler, operating system, virtual machine hypervisor etc., to obtain the runtime information. Instrumentation is the most commonly used method for realizing runtime information extraction, which inserts information collection function calls into the program at specific code point. Program instrumentation can be either static (labeled as “SSI” in Table 5 column 9) or dynamic (labeled as “SDI” in Table 5 column 9) (Fang et al., 2021; Park et al., 2019; Heinze et al., 2020). The former happens before the PUT runs, and can be performed at compile-time (Luo et al., 2019; Pham et al., 2020; Luo et al., 2020; Natella, 2022; Zou et al., 2021; Meng et al., 2021; Ba et al., 2022; Li et al., 2022a; Schumilo et al., 2022; Ren et al., 2021) or by directly rewriting the binary (Andronidis and Cadar, 2022). The latter happens while the PUT is running, leveraging tools such as DynamoRIO (Park et al., 2019) or Frida (Heinze et al., 2020) to inject hooking functions at specific code points to collect runtime information.
Externally-Observable-Behavior-Based methods are the most general class of methods, as it doesn’t rely on any support of the execution environment and can be used in a blackbox manner. There are various externally observable behaviors, such as the output of the program (labeled as “Resp” in Table 5 column 9) and the side-channel information such as power consumption and response time. The heuristics behind these observable behavior-based methods are that the differences in these behaviors can represent the PUT is under different states or having gone through different execution paths. Specifically, AFLNET (Pham et al., 2020) and Fieldfuzz (Bytes et al., 2022) identify different protocol states according to the status code in the response messages. Snipuzz (Feng et al., 2021) and FUME (Pearson et al., 2022) adopt a heuristic that different response messages mean different execution paths. Thus, they keep the input that can cause a different response as a seed for subsequent mutation testing, expecting to increase the coverage. Aafer et al. (Aafer et al., 2021) use the execution logs of the Android system as a feedback to refine the input generation grammars as the developers usually add log statements to indicate the detailed information about the input validation. By observing the side channel information such as the system status, power consumption, and response time, Flowfuzz (Sommer et al., 2017) determines whether the hardware switches have gone through different execution paths.
@@ 6. Bug Collector
To address the challenges in protocol fuzzing, existing works design both memory-safety bug oracles and non-memory-safety bug oracles according to various information sources.
| Year | Work | T | Target | Bug Detector | |
@@
| 2021 | ICSFuzzer(Fang et al., 2021) | ● | ICS | Crash Logs | - |
| 2021 | TCP-Fuzz(Zou et al., 2021) | ◐ | TCP | Fatal Signals | Inconsistency in Transmission & DE |
@@
| 2022 | PCFuzzer(Liu et al., 2022) | ● | PLC | Liveness Check & APB | - |
@@
| 2023 | LOKI(Ma et al., 2023) | ● | Blockchain | Fatal Signals | Incorrect State Transitions |
@@
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; T: Taxonomy; General: The fuzzer is not designed for a specific type of protocol; -: Not detectable; ESM: Error-Signaling Messages; APB: Abnormal Physical Behaviors; DE: Differences in Execution.
6.1. Memory-Safety Bug Oracles
Memory-Safety bugs include stack overflow, heap overflow, Use-After-Free (UAF) etc., which can lead to crashes. Existing technologies mostly observe their running status through different channels to determine whether a memory-safety bug is triggered. From the information source point of view, there are five types of the commonly used oracles for detecting memory-safety bugs, namely process fatal signals and sanitizers, crash logs and debug information, error-signaling messages, timeout and liveness checks and abnormal physical behaviors.
6.1.1. Fatal Signals and Sanitizers.
have been extensively utilized as a pivotal mechanism for bug detection in a plethora of contemporary studies (Somorovsky, 2016; Luo et al., 2019; Ruge et al., 2020; Luo et al., 2020; Natella, 2022; Zou et al., 2021; Ren et al., 2021; Ba et al., 2022; Poncelet et al., 2022). Predominantly, memory-safety bugs manifest through the overwriting of data or code pointers with invalid values, leading to critical process disruptions such as segmentation faults or process terminations, thereby generating fatal signals like SIGSEGV, SIGABRT, and others. Fuzzers can detect faults by checking whether the PUT process is dying from these signals. Addressing the subset of memory-safety bugs that do not immediately precipitate program crashes, fuzzers employ sanitizers. Sanitizers are bug detection tools, specifically engineered to identify and highlight unsafe or undesirable memory access patterns. Upon identifying such anomalies, sanitizers are designed to terminate the PUT, thereby signaling the presence of a potential bug (Serebryany et al., 2012; Song et al., 2019; Stepanov and Serebryany, 2015; Serebryany and Iskhodzhanov, 2009). Sanitizers can be enabled at compile-time(Somorovsky, 2016; Luo et al., 2019; Ruge et al., 2020; Luo et al., 2020; Natella, 2022; Zou et al., 2021; Ren et al., 2021; Ba et al., 2022) or enabled dynamically at runtime (Maier et al., 2020).
@@ 6.1. Memory-Safety Bug Oracles
Some works determine whether the PUTs are crashed by analyzing the system logs or debug information (Garbelini et al., 2020; Fang et al., 2021; Qu et al., 2021; Bytes et al., 2022; Garbelini et al., 2022a; Park et al., 2022; Wang and Wang, 2023). These system logs and debug information can be obtained by various channels. Specifically, ICSFuzzer leverages the Windows EventLog Service to detect crash events on Windows systems (Fang et al., 2021). Swentooth (Garbelini et al., 2020) and Braktooth (Garbelini et al., 2022a) propose to collect the startup messages or crash messages in the system logs leveraging the debug ports exposed by respective Bluetooth development boards. Startup message is an indicator of program crashes, as the Bluetooth devices have a watchdog program to reset the Bluetooth SoC when finding it is unresponsive. Wang et al. (Wang and Wang, 2023) leverage NLP technology to process the logs and detect unintended behavior of PUT. Differently, L2Fuzz (Park et al., 2022) and FieldFuzz (Bytes et al., 2022) identify crashes by checking whether a crash dump was generated.
@@
Many protocols use special responses or status codes to indicate internal errors, therefore can be used for bug detection. For example, L2Fuzz detects Bluetooth L2CAP vulnerabilities by checking whether the packet received contains an error-signaling message such as Connection Failed, Connection Aborted, Connection Reset, and Connection Refused (Park et al., 2022). These error messages indicate that the PUT may be crashed. OWFuzz uses the Deauth/Disassoc frames, management frames of the Wi-Fi protocol to terminate the communication, as an indicator of anomaly during fuzzing Wi-Fi protocol stacks (Cao, 2021).
@@
The abnormal physical behavior of the target device, e.g., startup sound, can also be used as a bug oracle. For example, when fuzzing the Bluetooth sound device, Braktooth uses the event of repeated startup sound as a bug oracle (Garbelini et al., 2022a). This is because Bluetooth devices will be restarted by the watchdog program when an error occurs and a startup sound will be played during the booting process. Differently, PCFuzzer (Liu et al., 2022) leverages an oscilloscope to collect the physical signal of the output module to monitor the target’s status.
@@
As most memory-safety bugs result in process crashes, fuzzers can detect these bugs by liveness detection. A common way to check the liveness of a target is to set a static timeout for a response (Gascon et al., 2015; Chen et al., 2018; Cao, 2021; Qu et al., 2021; Feng et al., 2021; Garbelini et al., 2022c; Song et al., 2022). If the response message from the target is not received by the time, it is determined that the target process is dead. This method is suitable for environments with limited debugging techniques, such as unable to obtain process signals or debug logs. However, setting a fixed timeout is a relatively coarse-grained method, which may introduce false positives due to network fluctuations or excessive load on the target. Some works propose several active liveness checks to mitigate the false positive issues. For example, Snipuzz (Feng et al., 2021) resends input sequences for multiple times to reduce false positives. IoTFuzzer (Chen et al., 2018), OWFuzz (Cao, 2021), and BadMesher (Qu et al., 2021) use heartbeat messages (e.g., ICMP messages) to infer the status of the PUT.

### v3

@@ 5.1.2. Execution Cost Reduction
1. Reset Strategy Selection. A reset strategy is mainly used to determine the appropriate time or execution point for performing a reset, which has a significant influence on the performance of fuzzing. An early reset may cause the target to terminate when it is still performing some tasks that may be vulnerable, and a late reset may affect the efficiency of the test. A common approach is to set a fixed time interval before resetting the execution. For example, AFLNET (Pham et al., 2020) allows the user to manually configure the time delays before restarting the PUT. However, this approach is relatively coarse-grained and it is hard to determine an appropriate time interval. In order to precisely control when to reset the execution, some works use program analysis to find the location that indicates the end of the execution of an iteration, and instrument the target program to terminate at these code locations (Andronidis and Cadar, 2022; Schumilo et al., 2022; Krupp et al., 2022; Qin et al., 2023). For example, AMPFuzz (Krupp et al., 2022) performs static analysis and injects termination calls into code branches that do not contain message-sending APIs. In addition, some works choose not to perform an execution reset after each fuzzing iteration for performance boost. For example, Charon (Zuo et al., 2022) leverage a program status inferring module to infer the time point at which the PUT has finished processing the packet, thereby detecting the coverage of specific inputs and avoiding the need to repeatedly restart the PUT to collect feedback. Similarly, SGFuzz (Ba et al., 2022) does not restart the PUT in every iteration. Instead, it performs a post-analysis to attribute the relationship between inputs and the target program’s behavior. Specifically, it collects all the inputs on which the PUT has been executed and minimizes the input list to a minimal message sequence that can trigger the bug.
@@
2. Execution State Reset. Execution state reset is responsible for resetting the context of the running PUT process to a specified state, including the data in registers and memory, etc. The existing execution state reset mechanisms can be divided into three categories, namely message-based reset, process restart, and snapshot & recovery.
@@
Message-based reset (labeled as “MR” in Table. 4 column 6) operates by sending a specific type of message that forces the PUT to terminate the ongoing session and revert to its initial state (Vanhoef, 2017; Fiterau-Brostean et al., 2020; Bytes et al., 2022). For instance, when fuzzing the Wi-Fi Access Point (AP), WiFuzz uses a deauthentication message to reset its state (Vanhoef, 2017). Message-based reset is easy-to-use, but it only supports a limited set of protocols as not every protocol is designed with a reset message. Furthermore, while it can reset the explicit protocol state of the PUT, it cannot reset the implicit state of the test target, such as global variables and memory that has been allocated but not freed.
@@
Another commonly used approach to reset the execution state is to kill the target process and restart (labeled as “ProcR” in Table. 4 column 6) (Chen et al., 2018; Luo et al., 2020; Fang et al., 2021). However, it is a relatively heavy operation for fuzzing, as the restart of the program involves multiple expensive pre-processing steps, such as loading the program into memory, dynamic linking, etc., resulting in inefficiencies.
@@
The snapshot & recovery mechanism has been integrated into fuzzing. This approach involves checkpointing PUT at a specific runtime state and then resetting it back to that checkpoint after each fuzzing iteration. This method effectively bypasses the repeated execution of resource-intensive initialization operations, thereby enhancing fuzzing efficiency. Protocol fuzzing, in particular, derives significant benefits from snapshot technology. Protocols are predominantly stateful, which implies that the input often comprises multiple prefix messages that guide the PUT to a designated state before introducing the crafted message. It is common for testcases to share the same prefix message sequences, especially when a specific state requires repeated exploration. Implementing snapshot technology in the protocol fuzzing process eliminates redundant executions associated with parsing these shared packet sequences, thereby markedly boosting fuzzing efficiency. The snapshot methodologies used in the current protocol fuzzing research can be broadly categorized into two types: process-level snapshots and virtual machine-level snapshots.
@@
Process-Level snapshot mechanisms (labeled as “PSR” in Table. 4 column 6) rely on system call capabilities provided by the operating system to realize their functionality. Generally, based on the APIs used, existing methods can be categorized into two types: fork-based and ptrace-based. Fork-based snapshot mechanisms are widely used in several well-known general-purpose fuzzers, including AFL(Zalewski, 2015). Specifically, AFL inserts a piece of fork-server code into the PUT program binary, which is executed before the function. Following a signal from the AFL fuzzing side, the fork-server generates a child process via the function, and this child process continues with the function. Since the fork-server has already loaded all kinds of resources, each child process only needs to execute the main function’s code, thereby bypassing the costly preprocessing steps and enhancing efficiency. This mechanism has been adopted by many protocol fuzzers for state resetting (Luo et al., 2019; Pham et al., 2020; Natella, 2022; Maier et al., 2020; Li et al., 2022a; Andronidis and Cadar, 2022; Qin et al., 2023; Fu et al., 2023). Furthermore, some works have extended the original fork-server mechanism in AFL to allow conditional multiple initializations at different code points, enabling the fuzzer to conveniently switch between various states of the protocol and thus boosting the fuzzing process (Chen et al., 2019; Andronidis and Cadar, 2022). Ptrace-based snapshot mechanisms, such as CRIU and DMTCP, leverage the debugging API to collect all the process context information and save it as image files (Li et al., 2022a). In the restoration process, these snapshot mechanisms read the dumped image files and recreate the process using syscalls such as or . Unlike the fork-based snapshot, which requires predetermined snapshot conditions (i.e., the location of the fork-server call) before execution, ptrace-based snapshots can checkpoint at any state during runtime.
@@
Virtual-Machine-Level snapshot mechanisms (labeled as “VMSR” in Table. 4 column 6) utilize the capabilities of virtual machine hypervisors to capture snapshots of the entire virtual machine at a specific time point, typically facilitated through a hypercall (Schumilo et al., 2022; Shi et al., 2019). When hypercalls are invoked, the program running within the virtual machine exits the VM context and transfers control to the hypervisor. Although the hypervisor-based approach is user-friendly, requiring no instrumentation, it is somewhat less efficient and more space-consuming because of its large granularity. To enhance the practicality of using virtual machine-level snapshots in protocol fuzzing, Nyx-net (Schumilo et al., 2022) implements an incremental snapshot approach to reduce the overhead associated with creating and removing snapshots. Specifically, Nyx-net establishes a root snapshot in pristine state, and each execution iteration commences from this root snapshot. In subsequent fuzzing iterations, Nyx-net generates incremental snapshots based on the root snapshot following the execution of an input message. Consequently, Nyx-net has a great performance boost on the testcases that share the same prefix message sequences.
@@
3. Execution Environment Reset. The reset of the execution environment primarily involves resetting the filesystem or database that may be affected by the PUT. Many fuzzers require users to provide a cleanup script to revert all changes (Pham et al., 2020; Qin et al., 2023; Fu et al., 2023; Li et al., 2022a; Natella, 2022), necessitating substantial manual effort to analyze the PUT’s potential impact on the external environment. To address this issue, Snapfuzz (Andronidis and Cadar, 2022) leverages a custom in-memory filesystem, where modifications are automatically discarded after the completion of a fuzzing iteration. Furthermore, the hypervisor-based snapshot mechanism. In addition, the hypervisor-based snapshot mechanism (VMSR in Table. 4 column 7) (Schumilo et al., 2022), which captures the state of the entire virtual machine, can reset both the execution state and the environment simultaneously.
@@ 5.2. Runtime Information Extraction
Software-Based methods leverage the capability of the software execution environment, e.g., compiler, operating system, virtual machine hypervisor, etc., to obtain the runtime information. Instrumentation is the most commonly used method for realizing runtime information extraction, which inserts information collection function calls into the program at specific code point. Program instrumentation can be static (labeled as “SSI” in Table 4 column 9) or dynamic (labeled as “SDI” in Table 4 column 9) (Fang et al., 2021; Park et al., 2019; Heinze et al., [n. d.]). The former happens before the PUT runs and can be performed at compile time (Luo et al., 2019; Pham et al., 2020; Luo et al., 2020; Natella, 2022; Zou et al., [n. d.]; Meng et al., 2021; Ba et al., 2022; Li et al., 2022a; Schumilo et al., 2022; Ren et al., 2021) or by directly rewriting the binary (Andronidis and Cadar, 2022). The latter happens while the PUT is running, leveraging tools such as DynamoRIO (Park et al., 2019) or Frida (Heinze et al., [n. d.]) to inject hooking functions at specific code points to collect runtime information.
Externally-Observable-Behavior-Based methods are the most general class of methods, as it does not rely on any support of the execution environment and can be used in a blackbox manner. There are various externally observable behaviors, such as the output of the program (labeled as “Resp” in Table 4 column 9) and the side-channel information such as power consumption and response time. The heuristics behind these observable behavior-based methods are that the differences in these behaviors can represent the PUT is under different states or having gone through different execution paths. Specifically, AFLNET (Pham et al., 2020) and Fieldfuzz (Bytes et al., 2022) identify different protocol states according to the status code in the response messages. Snipuzz (Feng et al., 2021) and FUME (Pearson et al., 2022) adopt the heuristic that different response messages mean different execution paths. Thus, they keep the input that can cause a different response as a seed for subsequent mutation testing, expecting to increase the coverage. Aafer et al. (Aafer et al., 2021) and Logos (Wu et al., 2024) use the execution logs as a feedback to refine the input generation grammars, as developers usually add log statements to indicate the detailed information about input validation. Observing side channel information such as system status, power consumption, and response time, Flowfuzz (Sommer et al., 2017) determines whether hardware switches have gone through different execution paths.
@@ 6. Bug Collector
To address the challenges in protocol fuzzing, existing work designs both memory-safety bug oracles and non-memory safety bug oracles according to various information sources, as shown in Fig. 7.
6.1. Memory-Safety Bug Oracles
6.1.1. Fatal Signals (Raised by Programs or Sanitizers).
have been extensively utilized as a pivotal mechanism for bug detection in a plethora of contemporary studies (Somorovsky, 2016; Luo et al., 2019; Ruge et al., [n. d.]; Luo et al., 2020; Natella, 2022; Zou et al., [n. d.]; Ren et al., 2021; Ba et al., 2022; Poncelet et al., 2022). Predominantly, memory-safety bugs manifest through the overwriting of data or code pointers with invalid values, leading to critical process disruptions such as segmentation faults or process terminations, thereby generating fatal signals like SIGSEGV, SIGABRT, and others. Fuzzers can detect faults by checking whether the PUT process is dying from these signals. Addressing the subset of memory-safety bugs that do not immediately precipitate program crashes, fuzzers employ sanitizers. Sanitizers are bug detection tools that are specifically engineered to identify and highlight unsafe or undesirable memory access patterns. Upon identification of such anomalies, sanitizers are designed to terminate PUT, thus signaling the presence of a potential bug (Serebryany et al., 2012; Song et al., 2019; Stepanov and Serebryany, 2015; Serebryany and Iskhodzhanov, 2009). Sanitizers can be enabled at compile-time(Somorovsky, 2016; Luo et al., 2019; Ruge et al., [n. d.]; Luo et al., 2020; Natella, 2022; Zou et al., [n. d.]; Ren et al., 2021; Ba et al., 2022) or enabled dynamically at runtime (Maier et al., 2020).
| Year | Work | Tax | Target | Bug Detector | |
@@ 6.1. Memory-Safety Bug Oracles
| 2021 | ICS3Fuzzer(Fang et al., 2021) | ● | ICS | Crash Logs | - |
| 2021 | TCP-Fuzz(Zou et al., [n. d.]) | ◐ | TCP | Fatal Signals | Inconsistency in Transmission & DE |
@@
| 2022 | PCFuzzer(Liu et al., 2022b) | ● | PLC | Liveness Check & APB | - |
@@
| 2023 | LOKI(Ma et al., 2023a) | ◐ | Blockchain | Fatal Signals | Incorrect State Transitions |
@@
○: Whitebox Fuzzer; ●: Blackbox Fuzzer; ◐: Greybox Fuzzer; Tax: Taxonomy; General: The fuzzer is not designed for a specific type of protocol; -: Not detectable; ESM: Error-Signaling Messages; APB: Abnormal Physical Behaviors; DE: Differences in Execution.
@@
Some works determine whether the PUTs are crashed by analyzing system logs or debug information (Garbelini et al., 2020; Fang et al., 2021; Qu et al., 2021; Bytes et al., 2022; Garbelini et al., 2022a; Park et al., 2022; Wang and Wang, 2023). These system logs and debug information can be obtained through various channels. Specifically, ICS3 Fuzzer uses the Windows EventLog Service to detect crash events on Windows systems (Fang et al., 2021). Swentooth (Garbelini et al., 2020) and Braktooth (Garbelini et al., 2022a) propose to collect startup messages or crash messages in the system logs using the debug ports exposed by respective Bluetooth development boards. The startup message is an indicator of program crashes, as Bluetooth devices have a watchdog program to reset the Bluetooth SoC when finding it is unresponsive. Wang et al. (Wang and Wang, 2023) leverage NLP technology to process the logs and detect unintended behavior of PUT. Unlike other sources, L2Fuzz (Park et al., 2022) and FieldFuzz (Bytes et al., 2022) identify crashes by checking whether a crash dump was generated.
@@
Many protocols use special responses or status codes to indicate internal errors, and therefore can be used for bug detection. For example, L2Fuzz detects Bluetooth L2CAP vulnerabilities by checking whether the received packet contains an error signaling message such as Connection Failed, Connection Aborted, Connection Reset, and Connection Refused (Park et al., 2022). These error messages indicate that the PUT may be crashed. OWFuzz uses the Deauth / Disassoc frames, the management frames of the Wi-Fi protocol to terminate the communication, as an indicator of anomaly during the fuzzing of the Wi-Fi protocol stacks (Cao, 2021).
@@
The abnormal physical behavior of the target device, e.g., startup sound, can also be used as a bug oracle. For example, when fuzzing the Bluetooth sound device, Braktooth uses the event of repeated startup sound as a bug oracle (Garbelini et al., 2022a). This is because Bluetooth devices will be restarted by the watchdog program when an error occurs, and a startup sound will be played during the booting process. Differently, PCFuzzer (Liu et al., 2022b) leverages an oscilloscope to collect the physical signal of the output module to monitor the target’s status.
@@
Timeout and liveness checks identify crashes or infinite loops by checking the target’s unresponsiveness. A common way to check the unresponsiveness of a target is to set a static timeout for a response (Gascon et al., 2015; Chen et al., 2018; Cao, 2021; Qu et al., 2021; Feng et al., 2021; Garbelini et al., 2022c; Song et al., 2022). If the response message from the target is not received by the time, it is determined that the target process is dead or enters an infinite loop. This method is suitable for environments with limited debugging techniques, such as unable to obtain process signals or debug logs. However, setting a fixed timeout is a relatively coarse-grained method, which may introduce false positives due to network fluctuations or excessive load on the target. Some works propose several active liveness checks to mitigate the false positive issues. For example, Snipuzz (Feng et al., 2021) resends input sequences for multiple times to reduce false positives. IoTFuzzer (Chen et al., 2018), OWFuzz (Cao, 2021), and BadMesher (Qu et al., 2021) use heartbeat messages (e.g., ICMP messages) to infer the status of the PUT.